Assume all future cash flows (FCFF, FCFE, and Dividends) will grow at a sonstant sustainable growth rate (g) in perpetuity; policy. The companys cost of debt is 2096 while the companys cost of equity is 30%; the company has an effective tax rate of 35%. Use the following information in the table below to help answer problems 26-27; FCFE Today (T 0) FCFF Today (T 0) Shareholder Equity Total Debt Total Assets Net Income in millions 800 750 400 600 1000 100 40 25 Shares Outstanding 26. An activist investor who wants to purchase all the companys shares would be willing to pay approximately Do not use the DDM. Use either FCFF or FCFE, whichever is appropriate a $165 per share b. $176 per share c $213 per share d. $230 per share e. $245 per share 27. A private equity group who wants to purchase all of the companys assets would be willing to pay approximately_Do not use the DDM. Use either FCFF or FCFE, whichever is appropriate a. $9,583M b. $10,222M с $10.518M d. $17,969M e. $19,167M
Answer #1

(26) Cost of Equity = 30 %, Net Income = 100 and Equity = 400, Return on Equity (ROE) = (100 / 400) = 0.25 or 25 %

Dividend = 40, Payout Ratio = Dividend / Net Income = 40 / 100 = 0.4

Retention Ratio = 1 - 0.4 = 0.6

Growth Rate = ROE x Retention Ratio = 25 x 0.6 = 15 %

FCFE = 800 (FCFE instead of FCFF used as one needs to determine the per share value which would require the equity value to be determined first)

Equity Value = [800 x 1.15] / [0.3 - 0.15] = $ 6133.33

Number of Shares Outstanding = 6133.33 / 25 = $ 245.33 ~ $ 245

Hence, the correct option is (e)

(27) In this case, the PE firm would need to know the total value of the firm which would mean that FCFF (and not just FCFE) should be used

Cost of Debt = 20 % and Tax Rate = 35 %

Debt Ratio = Debt / Assets = 600 / 100 = 0.6 and Equity Ratio =(1 - Debt Ratio) = (1 - 0.6) = 0.4

Weighetd Averaged Cost of Capital of the Firm (WACC) = 0.6 x 20 x (1-0.35) + 0.4 x 30 = 19.8%

Total Firm Value = [750 x 1.15] / [0.198 - 0.15] = $ 17968.75 ~ $ 17969

Hence, the correct option is (d)